OpenStudy (anonymous):

what does 100- (-45) equal????

OpenStudy (anonymous):

See : \(-(-)\) becomes \(+\)..

OpenStudy (anonymous):

ah i see @waterineyes

OpenStudy (anonymous):

So: \[100 \color{blue}{ - }(\color{blue}{-}45) \implies 100 + 45 = ??\]

OpenStudy (anonymous):

So, here you are to just add them..

OpenStudy (anonymous):

oh crap my question was supposed to be: what does 100+(-45)= uugggghhh whoops

OpenStudy (anonymous):

Then you will use: \(+(-)\) becomes \(-\).. So: \(100 - 45 = ??\)